# Quasiparticle interference and spectral function of the UTe2 superconductive surface band

### v1

@@ Acknowledgements.
A.C. thanks the CALMIP supercomputing center for the allocation of HPC numerical resources through project M23023. J.C.S.D. acknowledges support from the Royal Society under Award R64897. J.P.C., K.Z. and J.C.S.D. acknowledge support from Science Foundation Ireland under Award SFI 17/RP/5445. S.W. and J.C.S.D. acknowledge support from the European Research Council (ERC) under Award DLV-788932. Q.G., K.Z., J.P.C., S.W., and J.C.S.D. acknowledge support from the Moore Foundation’s EPiQS Initiative through Grant GBMF9457.
@@ A.1 Tight-binding parameters
Parameter DFT [22] (in eV) This work (in eV)
$\mu_{U}$0.35 -0.355
$\Delta_{U}$ 0.40 0.38
$t_{U}$ 0.15 0.17
$t^{\prime}_{U}$ 0.08 0.08
$t_{ch,U}$ 0.01 0.015
$t^{\prime}_{ch,U}$ 0.0 0.01
$t_{z,U}$0.03 -0.0375
$\mu_{Te}$1.80 -2.25
$\Delta_{Te}$1.50 -1.40
$t_{Te}$1.50 -1.50
$t_{ch,Te}$ 0.0 0.0
$t_{z,Te}$0.05 -0.05
$\delta$ 0.1 0.13
Table 3: List of parameters for the tight-binding model given in Ref. [22] (central column), and used in this work (right column).
@@ A.2 Bogoliubov-de-Gennes Hamiltonian
$\mathcal{H}_{\mathrm{BdG}}(\mathbf{k})=\begin{pmatrix}\mathcal{H}_{\mathrm{TB}}(\mathbf{k})\otimes\mathbb{1}_{4}&\Delta(\mathbf{k})\otimes\mathbb{1}_{2}\\ \Delta^{\dagger}(\mathbf{k})\otimes\mathbb{1}_{2}&-\mathcal{H}^{*}_{\mathrm{TB}}(-\mathbf{k})\otimes\mathbb{1}_{4}\\ \end{pmatrix}$ (31)
@@ A.3 Method to compute the surface Green function
For a body-centered orthorhombic lattice structure with the space group symmetry $Imm$, the reciprocal reciprocal lattice vectors are the following [41]
${\bf v}_{1}=\dfrac{2\pi}{ca}\begin{pmatrix}c\\ 0\\ a\end{pmatrix},\;{\bf v}_{2}=\dfrac{2\pi}{cb}\begin{pmatrix}0\\ -c\\ b\end{pmatrix},\;{\bf v}_{3}=\dfrac{2\pi}{ba}\begin{pmatrix}b\\ -a\\ 0\end{pmatrix}$
@@
$\mathcal{R}_{x}(\theta)=\begin{pmatrix}1&0&0\\ 0&\cos\theta&-\sin\theta\\ 0&\sin\theta&\cos\theta\end{pmatrix}$ (33)
@@ Appendix B Experimental constraints from symmetries
In this section, we present a detailed analysis of the symmetry constraints for chiral and non-chiral order parameters. As the QPI patterns depend on the orthogonality of eigenstates, certain peaks can vanish when the eigenstates are related by symmetries such as time-reversal symmetry $\mathcal{T}$. Thus implies that such considerations are important for guiding potential experimental studies. In Table 4 we summarize all the relevant symmetry constraints.
@@ B.1 Group Theory
Let’s consider an arbitrary point group $G$ such that $P$ is generated by the following crystalline symmetries $P=\langle g_{1},\cdots,g_{n}\rangle$ with $n$ the number of elements in the group. In the following, we only consider space groupe $P_{S}$ which are the direct product of translation symmetry and a point group $P$. The normal state Hamiltonian should transform as a representation $\rho$ of the space group. For $g\in P$,
$\mathcal{H}_{TB}(\mathbf{k})=\rho(g)\mathcal{H}_{TB}(g\mathbf{k})\rho(g)^{-1}$ (34)
The gap function $\Delta(\mathbf{k})$ defined by Eq. (9) is not invariant under this transformation and instead transform as [51]
$\Delta(\mathbf{k})=\rho(g)\Delta(g\mathbf{k})\rho^{T}(g)\Theta^{*}(g)$ (35)
where $\Theta(g)$ is a $1d$ representation of the symmetry group $P_{S}$. This means that $\theta$ is a scalar quantity and respect the group structure. However, we can define a set of matrices $\rho_{\mathcal{C}}(g)$ such that the BdG Hamiltonian defined in Eq. (31) transforms as
$\mathcal{H}_{\mathrm{BdG}}(\mathbf{k})=\rho_{\mathcal{C}}(g)\mathcal{H}_{\mathrm{BdG}}(g\mathbf{k})\rho_{\mathcal{C}}(g)^{-1}$ (36)
and $\rho_{\mathcal{C}}(g)$ forms a representation of $G$. It can be checked that $\rho_{\mathcal{C}}$ can be expressed in term of $\rho$ and $\Theta$ as [51]
$\rho_{\mathcal{C}}(g)=\begin{pmatrix}\rho(g)&0\\ 0&\rho^{*}(g)\Theta(g)\\ \end{pmatrix}$ (37)
@@
Finally $\mathcal{H}_{\mathrm{BdG}}(\mathbf{k})$ possesses an additional particle-hole symmetry $\mathcal{C}$, arising from the intrinsic redundancy of the particle-hole basis [52].
$\mathcal{H}_{\mathrm{BdG}}(\mathbf{k})=-\mathcal{C}\mathcal{H}_{\mathrm{BdG}}(-\mathbf{k})\mathcal{C}^{-1}$ (38)
The point-group of the normal state Hamiltonian discussed in this work is $D_{2h}$ [41].
$D_{2h}=\{E,m_{x},m_{y},m_{z},C_{2x},C_{2y},\mathcal{I}\}$ (39)
The normal state is also time-reversal symmetric. Time-reversal symmetry (TRS), denoted as $\mathcal{T}$, is an anti-unitary symmetry which is spinfull here such that $\mathcal{T}^{2}=-1$. The symmetry group including time-reversal symmetry $\mathcal{T}$ is called a magnetic group and is given by [41]
$D_{2h}^{\mathcal{T},II}=D_{2h}+\mathcal{T}D_{2h}$ (40)
@@
The symmetry group for non-chiral order parameters is also $D_{2h}^{\mathcal{T},II}$. However, chiral order parameters exhibit lower symmetry, as they are the sum of two different representations of $D_{2h}$, with a $\pi/2$ phase difference, causing them to transform differently under crystalline symmetries and breaking on-site time-reversal symmetry $\mathcal{T}$. Their symmetry group corresponds to a type-III magnetic point group [41]. The magnetic point group for $A_{u}+iB_{1u}$ and $B_{2u}+iB_{3u}$ is
$D_{2h}^{\mathcal{T},III}=\langle m_{z},C_{2z},\mathcal{I},\mathcal{T}m_{y},\mathcal{T}C_{2y},\mathcal{T}m_{x},\mathcal{T}C_{2}\rangle$ (41)
For $A_{u}+iB_{2u}$ and $B_{1u}+iB_{3u}$ it read
$D^{\mathcal{T},III}_{2h}=\langle m_{y},C_{2y},\mathcal{I},\mathcal{T}m_{z},\mathcal{T}C_{2z},\mathcal{T}m_{x},\mathcal{T}C_{2x}\rangle$ (42)
@@
$D^{\mathcal{T},III}_{2h}=\langle m_{x},C_{2x},\mathcal{I},\mathcal{T}m_{y},\mathcal{T}C_{2y},\mathcal{T}m_{z},\mathcal{T}C_{2z}\rangle$ (43)
@@ B.2 Symmetry constraints on Green’s function
Eq. (36) implies that the BdG Green’s function defined in Eq. (14) transforms under a point-group transformation $g$ as
$G_{\mathrm{BdG}}(E,\mathbf{k})=\rho(g)^{-1}G_{\mathrm{BdG}}(E,g\mathbf{k})\rho(g)$ (44)
Eq. (38) implies that under particle-hole symmetry $\mathcal{C}$, the Green’s function transforms as follows
$G_{\mathrm{BdG}}(E,\mathbf{k})=\mathcal{C}^{-1}G_{\mathrm{BdG}}(-E,-\mathbf{k})\mathcal{C}$ (45)
As discussed in Sec. II.3, the surface spectrum consists of two distinct contributions: one purely from the bulk which is independent of the edge impurity denoted as $G_{b}$, and another arising solely from edge effects named $G_{st}$. Their expressions are given respectively in Eq. (II.3) and Eq. (II.3) such that the physical surface Green’s function is the sum of these two contributions.
In the following, we considered only symmorphic crystalline symmetries, which means $\rho(g)$ is $k$-independent. The model form a representation of the space group $P_{S}$ which is then the semi-direct product of the $3D$ translation group $T_{3}$ and the point group $P=D_{2h}$, $P_{S}=D_{2h}\rtimes T_{3}$. $T_{3}$ is composed of the following elements $t$, $t=n_{1}t_{1}+n_{2}t_{2}+n_{3}t_{3}$ where $n_{i}\in\mathds{N}$ and $t_{i}$ are elementary translations that join a point with its nearest neighbors. If we write a group element as $g^{a/b}_{s}=\{g^{a/b}_{i}|t^{a/b}\}$ with $g^{a/b}\in P$ and $t^{a/b}\in T_{3}$, then the semi-product $\rtimes$ means
$\{g_{i}^{a}|t^{a}\}\{g_{i}^{b}|t^{b}\}=\{g_{i}^{a}g_{i}^{b}|t^{a}+g_{i}^{a}t_{b}\}$ (46)
@@
[…]
@@
[…]
Then for each pairing symmetry, the study of Eq. (47) and Eq. (48) with their point group symmetry is constraining the surface spin-resolved and non-spin-resolved spectral function.
@@ B.4 Transformation of $G_{i}$
For the symmetry transformation of $G_{i}$, two distinct cases must be considered for each order parameter. This distinction arises from the particle-hole transformation, under which $G_{\mathrm{BdG}}(E,\mathbf{k})$ transforms as $G_{\mathrm{BdG}}(-E,-\mathbf{k})$. When $E=0$, $G_{\mathrm{BdG}}(E=0,\mathbf{k})\rightarrow G_{\mathrm{BdG}}(E=0,-\mathbf{k})$ so $G_{\mathrm{BdG}}(E=0,\mathbf{k})$ is constrained by particle-hole symmetry. In contrast, for $E\neq 0$, $G_{\mathrm{BdG}}(E,\mathbf{k})\rightarrow G_{\mathrm{BdG}}(-E,-\mathbf{k})$ so there is no direct constraint on $G_{\mathrm{BdG}}(E,\mathbf{k})$.

### v2

@@ Acknowledgements.
A.C. thanks the CALMIP supercomputing center for the allocation of HPC numerical resources through Project M23023 supported by a French government grant managed by the Agence Nationale de la Recherche under the Investissements d’avenir program (ANR-21-ESRE-0051). J.C.S.D. acknowledges support from the Royal Society under Award R64897. J.P.C., K.Z. and J.C.S.D. acknowledge support from Science Foundation Ireland under Award SFI 17/RP/5445. S.W. and J.C.S.D. acknowledge support from the European Research Council (ERC) under Award DLV-788932. Q.G., K.Z., J.P.C., S.W., and J.C.S.D. acknowledge support from the Moore Foundation’s EPiQS Initiative through Grant GBMF9457. K.Z. acknowledges funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie Actions Postdoctoral Global Fellowship “QUASAR”, grant agreement No. 101203931.
@@ A.1 Tight-binding parameters
To describe the band structure of UTe2, we use a 4-orbital tight-binding model introduced by Theuss et al. [theuss2024single], and parameter values modified compared with those obtained from DFT or QO calculations. The parameters entering in Eq. (1) are given in the last column of Tab. 3. It leads to an area for the electron and hole pockets of $\mathcal{A}_{e}\approx\mathcal{A}_{h}\approx 34.9$ nm-2 at $k_{c}=0$, a value which is close to the value ($\mathcal{A}_{e}\approx\mathcal{A}_{h}\approx\ 33.6$ nm-2) measured in recent experiments [eaton2024quasi, weinberger2024quantum]. As shown in Fig. 19, the topology of the band structure close to energy equal to zero is qualitatively similar, with two bands crossing the Fermi level, one with a maximum close to the $\Sigma$ point, and one with a minimum close to the $\Delta$ point, giving rise respectively to the electron and hole pockets consistent with the quantum oscillation measurements [eaton2024quasi, weinberger2024quantum]. We also note that, for our parameter set, the band exhibits a flattening near zero energy along the $\Sigma X$ direction close to the $X$ point, in a manner quite similar to what is observed for the QO parameter set of Ref. [theuss2024single]. The exact form of the bands far away from the Fermi level does not affect the physics since we focus on SC gaps of the order of sub-mV.
@@
Parameter DFT [theuss2024single] QO [theuss2024single] This work
$\mu_{U}$ 0.35 0.35 -0.355
$\Delta_{U}$ 0.4 0.05 0.38
$t_{U}$ 0.15 0.1 0.17
$t^{\prime}_{U}$ 0.08 0.08 0.08
$t_{ch,U}$ 0.01 0.01 0.015
$t^{\prime}_{ch,U}$ 0 0 0.01
$t_{z,U}$ -0.03 0.04 -0.0375
$\mu_{Te}$ -1.8 -1.8 -2.25
$\Delta_{Te}$ -1.5 -1.5 -1.4
$t_{Te}$ -1.5 -1.5 -1.5
$t_{ch,Te}$ 0 -0.03 0
$t_{z,Te}$ -0.05 -0.5 -0.05
$\delta$ 0.09 0.1 0.13
Table 3: List of parameters in eV for the tight-binding model given in Ref. [theuss2024single] (central columns), and used in this work (last column).
@@ A.2 Bogoliubov-de-Gennes Hamiltonian
$\mathcal{H}_{\mathrm{BdG}}(\mathbf{k})=\begin{pmatrix}\mathcal{H}_{\mathrm{TB}}(\mathbf{k})\otimes\mathbb{1}_{2}&\Delta(\mathbf{k})\otimes\mathbb{1}_{4}\\ \Delta^{\dagger}(\mathbf{k})\otimes\mathbb{1}_{4}&-\mathcal{H}^{*}_{\mathrm{TB}}(-\mathbf{k})\otimes\mathbb{1}_{2}\\ \end{pmatrix}$, (31)
@@ A.3 Method to compute the surface Green function
For a body-centered orthorhombic lattice structure with the space group symmetry $Imm$, the reciprocal reciprocal lattice vectors are the following [bradley2009mathematical]
${\bf v}_{1}=\dfrac{2\pi}{ca}\begin{pmatrix}c\\ 0\\ a\end{pmatrix},\;{\bf v}_{2}=\dfrac{2\pi}{cb}\begin{pmatrix}0\\ -c\\ b\end{pmatrix},\;{\bf v}_{3}=\dfrac{2\pi}{ba}\begin{pmatrix}b\\ -a\\ 0\end{pmatrix}~$.
@@
$\mathcal{R}_{x}(\theta)=\begin{pmatrix}1&0&0\\ 0&\cos\theta&-\sin\theta\\ 0&\sin\theta&\cos\theta\end{pmatrix}~$, (33)
@@
A.4 DOS with additional d-vector contributions
For the sake of completeness, we calculate the DOS when adding allowed symmetry contributions to the d-vector such as
$\mathbf{d}_{B_{2u}}(\mathbf{k})=\begin{pmatrix}\Delta_{0}\sin(ck_{z})\\ C_{0}\sin(ak_{x})\sin(bk_{y})\sin(ck_{z})\\ \Delta_{0}\sin(ak_{x})\end{pmatrix}~$, (34)
and
$\mathbf{d}_{B_{3u}}(\mathbf{k})=\begin{pmatrix}C_{0}\sin(ak_{x})\sin(bk_{y})\sin(ck_{z})\\ \Delta_{0}\sin(ck_{z})\\ \Delta_{0}\sin(bk_{y})\end{pmatrix}~$. (35)
The result is displayed in Fig. 20. It shows that the peak in the B3u surface DOS obtained at zero energy when $C_{0}=0$ splits in two side-peaks when $C_{0}=\Delta_{0}$.
Figure 20: Bulk and surface DOS for B2u and B3u pairings with $C_{0}=\Delta_{0}=0.3$ meV and $\eta=0.1$ meV.
@@ Appendix B Experimental constraints from symmetries
In this section, we present a detailed analysis of the symmetry constraints for chiral and non-chiral order parameters. As the QPI patterns depend on the orthogonality of eigenstates, certain peaks can vanish when the eigenstates are related by symmetries such as time-reversal symmetry $\mathcal{T}$. Thus implies that such considerations are important for guiding potential experimental studies. In Tables 4 and 5 we summarize all the relevant symmetry constraints.
@@ B.1 Group Theory
Let us consider an arbitrary point group $G$ such that $P$ is generated by the following crystalline symmetries $P=\langle g_{1},\cdots,g_{n}\rangle$ with $n$ the number of elements in the group. In the following, we only consider the space group $P_{S}$ which is the direct product of the translation symmetry and a point group $P$. The normal state Hamiltonian should transform as a representation $\rho$ of the space group. For $g\in P$,
$\mathcal{H}_{TB}(\mathbf{k})=\rho(g)\mathcal{H}_{TB}(g\mathbf{k})\rho(g)^{-1}~.$ (36)
The gap function $\Delta(\mathbf{k})$ defined by Eq. (9) is not invariant under this transformation and instead transforms as [geier2020symmetry]
$\Delta(\mathbf{k})=\rho(g)\Delta(g\mathbf{k})\rho^{T}(g)\Theta^{*}(g)~,$ (37)
where $\Theta(g)$ is a $1D$ representation of the symmetry group $P_{S}$. This means that $\Theta$ is a scalar quantity and respects the group structure. However, we can define a set of matrices $\rho_{\mathcal{C}}(g)$ such that the BdG Hamiltonian defined in Eq. (31) transforms as
$\mathcal{H}_{\mathrm{BdG}}(\mathbf{k})=\rho_{\mathcal{C}}(g)\mathcal{H}_{\mathrm{BdG}}(g\mathbf{k})\rho_{\mathcal{C}}(g)^{-1}~,$ (38)
and $\rho_{\mathcal{C}}(g)$ forms a representation of $G$. It can be checked that $\rho_{\mathcal{C}}$ can be expressed in terms of $\rho$ and $\Theta$ as [geier2020symmetry]
$\rho_{\mathcal{C}}(g)=\begin{pmatrix}\rho(g)&0\\ 0&\rho^{*}(g)\Theta(g)\\ \end{pmatrix}~$. (39)
@@
Finally $\mathcal{H}_{\mathrm{BdG}}(\mathbf{k})$ possesses an additional particle-hole symmetry $\mathcal{C}$, arising from the intrinsic redundancy of the particle-hole basis [chiu2016classification].
$\mathcal{H}_{\mathrm{BdG}}(\mathbf{k})=-\mathcal{C}\mathcal{H}_{\mathrm{BdG}}(-\mathbf{k})\mathcal{C}^{-1}~.$ (40)
The point-group of the normal state Hamiltonian discussed in this work is $D_{2h}$ [bradley2009mathematical].
$D_{2h}=\{E,m_{x},m_{y},m_{z},C_{2x},C_{2y},\mathcal{I}\}~.$ (41)
The normal state is also time-reversal symmetric. Time-reversal symmetry (TRS), denoted as $\mathcal{T}$, is an anti-unitary symmetry which is spinfull here such that $\mathcal{T}^{2}=-1$. The symmetry group including time-reversal symmetry $\mathcal{T}$ is called a magnetic group and is given by [bradley2009mathematical]
$D_{2h}^{\mathcal{T},II}=D_{2h}+\mathcal{T}D_{2h}~,$ (42)
@@
The symmetry group for non-chiral order parameters is also $D_{2h}^{\mathcal{T},II}$. However, chiral order parameters exhibit lower symmetry, as they are the sum of two different representations of $D_{2h}$, with a $\pi/2$ phase difference, causing them to transform differently under crystalline symmetries and breaking on-site time-reversal symmetry $\mathcal{T}$. Their symmetry group corresponds to a type-III magnetic point group [bradley2009mathematical]. The magnetic point group for $A_{u}+iB_{1u}$ and $B_{2u}+iB_{3u}$ is
$D_{2h}^{\mathcal{T},III}=\langle m_{z},C_{2z},\mathcal{I},\mathcal{T}m_{y},\mathcal{T}C_{2y},\mathcal{T}m_{x},\mathcal{T}C_{2}\rangle~.$ (43)
For $A_{u}+iB_{2u}$ and $B_{1u}+iB_{3u}$ it reads
$D^{\mathcal{T},III}_{2h}=\langle m_{y},C_{2y},\mathcal{I},\mathcal{T}m_{z},\mathcal{T}C_{2z},\mathcal{T}m_{x},\mathcal{T}C_{2x}\rangle~.$ (44)
@@
$D^{\mathcal{T},III}_{2h}=\langle m_{x},C_{2x},\mathcal{I},\mathcal{T}m_{y},\mathcal{T}C_{2y},\mathcal{T}m_{z},\mathcal{T}C_{2z}\rangle~.$ (45)
@@ B.2 Symmetry constraints on Green’s function
Equation (38) implies that the BdG Green’s function defined in Eq. (14) transforms under a point-group transformation $g$ as
$G_{\mathrm{BdG}}(E,\mathbf{k})=\rho(g)^{-1}G_{\mathrm{BdG}}(E,g\mathbf{k})\rho(g)~.$ (46)
Equation (40) implies that under particle-hole symmetry $\mathcal{C}$, the Green’s function transforms as follows
$G_{\mathrm{BdG}}(E,\mathbf{k})=\mathcal{C}^{-1}G_{\mathrm{BdG}}(-E,-\mathbf{k})\mathcal{C}~.$ (47)
As discussed in Sec. II.3, the surface spectrum consists of two distinct contributions: one purely from the bulk which is independent of the edge impurity denoted $G_{b}$, and another arising solely from edge effects named $G_{i}$. Their expressions are given respectively in Eqs. (II.3) and (II.3) such that the physical surface Green’s function is the sum of these two contributions.
In the following, we consider only symmorphic crystalline symmetries, which means $\rho(g)$ is $k$-independent. The model form a representation of the space group $P_{S}$ which is then the semi-direct product of the $3D$ translation group $T_{3}$ and the point group $P=D_{2h}$, $P_{S}=D_{2h}\rtimes T_{3}$. $T_{3}$ is composed of the elements $t$, $t=n_{1}t_{1}+n_{2}t_{2}+n_{3}t_{3}$ where $n_{i}\in\mathds{N}$ and $t_{i}$ are elementary translations that join a point with its nearest neighbors. If we write a group element as $g^{a/b}_{s}=\{g^{a/b}_{i}|t^{a/b}\}$ with $g^{a/b}\in P$ and $t^{a/b}\in T_{3}$, then the semi-product $\rtimes$ means
$\{g_{i}^{a}|t^{a}\}\{g_{i}^{b}|t^{b}\}=\{g_{i}^{a}g_{i}^{b}|t^{a}+g_{i}^{a}t_{b}\}~.$ (48)
@@
$G(E,k_{x},k_{\#},x)=\rho^{-1}_{\mathcal{C}}(g)$
$\times\left(\int dk_{\perp}e^{ixk_{\perp}}G_{\mathrm{BdG}}(E,g\mathbf{k})\right)\rho_{\mathcal{C}}(g)~.$ (49)
@@
$\times\left(\int dk_{\perp}e^{-ixk_{\perp}}G_{\mathrm{BdG}}(E,-g_{b}\mathbf{k})\right)\rho^{A}_{\mathcal{C}}(\mathcal{T}g_{b})~.$ (50)
Then, for each pairing symmetry, the study of Eqs. (B.2) and (50) with their point group symmetry is constraining the surface spin-resolved and non-spin-resolved spectral function.
@@ B.4 Transformation of $G_{i}$
For the symmetry transformation of $G_{i}$, two distinct cases must be considered for each order parameter. This distinction arises from the particle-hole transformation, under which $G_{\mathrm{BdG}}(E,\mathbf{k})$ transforms into $G_{\mathrm{BdG}}(-E,-\mathbf{k})$. When $E=0$, $G_{\mathrm{BdG}}(E=0,\mathbf{k})\rightarrow G_{\mathrm{BdG}}(E=0,-\mathbf{k})$ so $G_{\mathrm{BdG}}(E=0,\mathbf{k})$ is constrained by particle-hole symmetry. In contrast, for $E\neq 0$, $G_{\mathrm{BdG}}(E,\mathbf{k})\rightarrow G_{\mathrm{BdG}}(-E,-\mathbf{k})$ so there is no direct constraint on $G_{\mathrm{BdG}}(E,\mathbf{k})$.